OpenStudy (anonymous):

Solve each equation. ½ + 3/t = 5/8

OpenStudy (anonymous):

t=24

OpenStudy (anonymous):

how'd you get 24?

OpenStudy (anonymous):

3/t=5/8-1/2 to get common denominator multiply 1/2 by 4/4

OpenStudy (anonymous):

you get 4/8

OpenStudy (anonymous):

5/8-4/8=1/8

OpenStudy (anonymous):

3/t=1/8

OpenStudy (anonymous):

multiply both sides by t 3=(1/8)*t and finally t=8*3=24

OpenStudy (anonymous):

oh ok thanks can you do 2 more?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3/m-4 + 1/3(m-4) = 6/m

OpenStudy (anonymous):

1/3 (m+9/m-16) = 6/m m =8+sqrt(73) and m =-8+sqrt(73)

OpenStudy (anonymous):

sorry m= 8-sqrt(73)

OpenStudy (anonymous):

so it's 9 right?

OpenStudy (anonymous):

there are two answers

OpenStudy (anonymous):

Oh okay, last one: 2c/c-4 – 2 = 4/c+5

OpenStudy (anonymous):

4/c+9 = 0 c=-4/9

OpenStudy (anonymous):

thanks & okay I have one more and after this, that's it: 5/2x – 3 = 7/3x

OpenStudy (anonymous):

x/6-3 = 0 x=18

OpenStudy (anonymous):

are you sure? because in the back of my friends book, 5/2x – 3 = 7/3x = -21 and for the previous question, 2c/c-4 – 2 = 4/c+5 = -14

OpenStudy (anonymous):

5/2x – 3 = 7/3x the answer was -21 and 2c/c-4 – 2 = 4/c+5 the answer was -14